A gas has a density of 5.89 g/L at a pressure of 1 atm and at a temperature of 20 degrees Celsius. What is the volume in liters of 5.89 kg of this gas under head conditions? Give your answer in scientific notation without units. You may disregard significant figures.
1
Expert's answer
2020-09-05T14:41:19-0400

5.89 kg = 5.89*103 g

If conditions remain constant, the volume will be:

5.89*103g * 1L / 5.89g = 1*103 L

Answer: 1*103